what is the range of the function rule y=5x-2 for the domain {0.5, 11}
Answer by Mr.positive(13)

What is the range of the function rule y=5x-2 for the domain {0.5, 11}
The values of the domain are the values of x.
The values of range are the values of y.
Take x = 0.5 and substitute it in the equation:
y = 5x – 2  y = 5(0.5) – 2  y = 2.5 – 2 = 0.5
Now take x = 11 and substitute it in the equation:
y = 5x – 2  y = 5(11) – 2  y = 55 – 2 = 49
Remember, the values of the range are the values of y that we obtained above.
Therefore, the values of the range are {0.5,49).
mathteacher2010@aol.com